OpenStudy (anonymous):

Write the equation of the line in slope intercept: form with a slope of 1 and a y intercept of -1

OpenStudy (australopithecus):

EQUATION OF THE LINE. MEMORIZE!!!! y = mx + b m = slope b = y-intercept

OpenStudy (anonymous):

but is that the answer...?

OpenStudy (anonymous):

lol for real? you just plug in what you have into the equation.

OpenStudy (anonymous):

lol yeah man for real

OpenStudy (anonymous):

y=1x-1

OpenStudy (anonymous):

well m-1=y

OpenStudy (anonymous):

mhm...

OpenStudy (australopithecus):

the answer is y = (1)x + (-1) y = x - 1

OpenStudy (anonymous):

thank you.

OpenStudy (australopithecus):

Remember A negative * A positive = A Negative hence why +(-1) = -1 or 1(-1) = -1

OpenStudy (anonymous):

Yeah i understand it now, thanks for your help.